can you show me how to do this: dirct variation equations
X 1 2 3 4 6
Y 5 10 15 20 30
For the formula of y=ax

Answer by checkley77(12844)

y=ax
1=a*5
a=1/5 answer.
Proof:
y=x/5
1=5/5
1=1
y=x/5
6=30/5
6=6